### v2

@@ 5.1 Local linear convergence
$\bar{H}:=\nabla^{2}f(\bar{x}).$ (29)
Also, for $u\in\mathbb{R}^{d}$ small enough, we have by setting $y=x_{k}+P_{k}^{\top}u$ in (28), that for $k$ large enough such that $x_{k}+P_{k}^{\top}u\in\hat{B}$,
$|f(x_{k}+P_{k}^{\top}u)-f(\bar{x})-\frac{1}{2}(x_{k}-\bar{x})^{\top}\bar{H}(x_{k}-\bar{x})-\frac{1}{2}u^{\top}P_{k}\bar{H}P_{k}^{\top}u-(P_{k}\bar{H}(x_{k}-\bar{x}))^{\top}u|$ (30)
@@
holds. Let $g(u)=\frac{1}{2}u^{\top}P_{k}\bar{H}P_{k}^{\top}u+(P_{k}\bar{H}(x_{k}-\bar{x}))^{\top}u$. By a reasoning similar to that of Proposition 5.1, $g$ is strongly convex with probability $1-6e^{-s}$ and hence is minimized at
$u^{*}=-(P_{k}\bar{H}P_{k}^{\top})^{-1}P_{k}\bar{H}(x_{k}-\bar{x}).$ (31)
Notice that as $k$ tends to infinity $\|u^{*}\|$ tends to 00, hence for $k$ large enough we have $x_{k}+P_{k}^{\top}u^{*}\in\hat{B}$ and $u^{*}\in B_{0}$. Plugging (31) in (30) yields
@@
where $\Pi=\bar{H}P_{k}^{\top}(P_{k}\bar{H}P_{k}^{\top})^{-1}P_{k}\bar{H}$. Using (28), we further obtain
$f(x_{k})-f(x_{k}+P_{k}^{\top}u^{*})\geq\frac{1}{2}(x_{k}-\bar{x})^{\top}\Pi(x_{k}-\bar{x})-L_{H}(\|x_{k}-\bar{x}+P_{k}^{\top}u^{*}\|^{3}+\|x_{k}-\bar{x}\|^{3}).$ (32)
@@
$=\frac{\lambda_{0}}{2\lambda_{\max}(P_{k}\bar{H}P_{k})}(x_{k}-\bar{x})^{\top}\bar{H}(x_{k}-\bar{x}).$ (33)
where the second inequality holds, with probability at least $1-2\exp(-\frac{\mathcal{C}_{0}}{4}s)$ (by Lemma 2.2, as $\lambda_{0}$ is the smallest non-zero eigenvalue of $\bar{H}$. The second equality holds as $\sigma_{\max}(P_{k}\bar{H}^{1/2})^{2}=\lambda_{\max}(P_{k}\bar{H}_{k}P_{k})$. We have therefore proved that
$(\bar{H}^{1/2}(x_{k}-\bar{x}))^{\top}\bar{\Pi}(\bar{H}^{1/2}(x_{k}-\bar{x}))\geq\frac{\lambda_{0}}{2\lambda_{\max}(P_{k}\bar{H}P_{k})}(x_{k}-\bar{x})^{\top}\bar{H}(x_{k}-\bar{x}).$ (34)
Hence, by (32), we have
@@
$-L_{H}(\|x_{k}-\bar{x}+P_{k}^{\top}u^{*}\|^{3}+\|x_{k}-\bar{x}\|^{3}).$ (35)
From (31), we have that $\|x_{k}-\bar{x}+P_{k}^{\top}u^{*}\|=\|(I_{n}-P_{k}^{\top}(P_{k}\bar{H}P_{k}^{\top})^{-1}P_{k}\bar{H})(x_{k}-\bar{x})\|$. Hence
$\|x_{k}-\bar{x}+P_{k}^{\top}u^{*}\|\leq\|I_{n}-P_{k}^{\top}(P_{k}\bar{H}P_{k}^{\top})^{-1}P_{k}\bar{H}\|\|x_{k}-\bar{x}\|.$ (36)
@@
$P_{k}\bar{H}P_{k}^{\top}\succeq\frac{(1-\varepsilon_{0})^{2}n}{2s}\sigma\bar{\lambda}I_{s}.$
Hence, we deduce from (37) that
$\|I_{n}-P_{k}^{\top}(P_{k}\bar{H}P_{k}^{\top})^{-1}P_{k}\bar{H}\|\leq\frac{\|P_{k}^{\top}\|^{2}\|\bar{H}\|}{\frac{(1-\varepsilon_{0})^{2}n}{2s}\sigma\bar{\lambda}}.$ (38)
Therefore, we deduce by (5.6), (36) and (38) for $\beta_{1}>0$ suitably chosen, we have
$f(x_{k})-f(x_{k}+P_{k}^{\top}u^{*})\geq\frac{\lambda_{0}}{4\lambda_{\max}(P_{k}\bar{H}P_{k})}(x_{k}-\bar{x})^{\top}\bar{H}(x_{k}-\bar{x})-\beta_{1}\|x_{k}-\bar{x}\|^{3}.$ (39)
By taking $y=x_{k}$ in (28), we have that
@@
Since $\|x_{k}-\bar{x}\|$ tends to 00 as $k$ tends to infinity and $\rho<3$, we have that for $k$ large enough
@@
where the first inequality holds as, by (31), $u^{*}\in B_{0}$ for $k$ large enough. The probability bound in the statement of the theorem is obtained by using (4): in the whole proof we only use Lemma 2.1 with $\varepsilon=\frac{1}{2}$, which holds with probability at least $1-2\exp(-\frac{\mathcal{C}_{0}}{4}s)$, and the inequalities (26) which hold with probability at least $1-6\exp(-s)$. We also factorize the expression, using that $1-2\exp(-\frac{\mathcal{C}_{0}}{4}s)>1-6\exp(-\frac{\mathcal{C}_{0}}{4}s)$. We end the proof by noticing that $\lambda_{\max}(P_{k}\bar{H}P_{k})\leq\lambda_{\max}(\bar{H})\sigma_{\max}(P_{k})^{2}$, hence by the first equation of (26)
$\lambda_{\max}(P_{k}\bar{H}P_{k})\leq\lambda_{\max}(\bar{H})\left(\sqrt{\frac{n}{s}}+\mathcal{C}\right)^{2}.$ (40)
@@
Assume that Assumptions 5, 5.1, and 5.1 hold. There exist $10<\kappa<$ 1, $k_{0}\in\mathbb{N}$, and $n_{0}\in\mathbb{N}$ such that if $n\geq n_{0}$ and $k\geq k_{0}$, then
@@
with $d_{k}=P_{k}^{\top}u_{k}$ and the update rule $t_{k}\leftarrow\beta t_{k}$ for $10<\alpha<$ 1 and $10<\beta<$ 1. We recall that
$u_{k}=-(P_{k}H_{k}P_{k}^{\top}+\eta_{k}I_{s})^{-1}P_{k}g_{k},$ (41)
@@
$f(x_{k}+t_{k}P_{k}^{\top}u_{k})=f(x_{k})+t_{k}(P_{k}g_{k})^{\top}u_{k}+\frac{t_{k}^{2}}{2}u_{k}^{\top}P_{k}\nabla^{2}f(x_{k}^{*})P_{k}u_{k}.$ (42)
Notice that $\nabla^{2}f$ is Lipschitz continuous (by Assumption 5). Furthermore, by Proposition 5.1, for $k$ large enough, $P_{k}{H}_{k}P_{k}$ is positive definite with probability at least $1-6\exp(-s)$ as the sequence $\{x_{k}\}$ converges to $\bar{x}$. Hence, we deduce that there exists $n_{0}$ such that for $n\geq n_{0}$ and for $k$ large enough
@@
holds with probability at least $1-6(\exp(-s)+\exp(-\frac{\mathcal{C}_{0}}{4}s))$. By (42), we deduce that for $k$ large enough and $n\geq n_{0}$:
$f(x_{k}+t_{k}P_{k}^{\top}u_{k})$ $\leq f(x_{k})+t_{k}(P_{k}g_{k})^{\top}u_{k}+2\frac{t_{k}^{2}}{2}u_{k}^{\top}P_{k}{H}_{k}P_{k}u_{k}$
$\leq f(x_{k})+t_{k}(P_{k}g_{k})^{\top}u_{k}+{t_{k}^{2}}u_{k}^{\top}(P_{k}{H}_{k}P_{k}+\eta_{k}I_{s})u_{k},$
where that second inequality holds as $\eta_{k}\geq 0$. Let
$\mu_{k}^{2}=-g_{k}^{\top}d_{k}=(P_{k}g_{k})^{\top}(P_{k}H_{k}P_{k}^{\top}+\eta_{k}I_{s})^{-1}(P_{k}g_{k}).$ (43)
Since $(P_{k}g_{k})^{\top}u_{k}=g_{k}^{\top}(P_{k}^{\top}u_{k})=-\mu_{k}^{2}$, and by definition of $u_{k}$, (41), we can write
$f(x_{k}+t_{k}P_{k}^{\top}u_{k})\leq f(x_{k})-t_{k}\mu_{k}^{2}+{t_{k}^{2}}u_{k}^{\top}(P_{k}{H}_{k}P_{k}+\eta_{k}I_{s})u_{k}=f(x_{k})-t_{k}\mu_{k}^{2}+{t_{k}^{2}}\mu_{k}^{2}.$ (44)
@@
for such $t_{k}$. Therefore, the backtracking line search stops with some $t_{k}\geq 1-\alpha$. Thus, when the line search exits we have
$f(x_{k+1})\leq f(x_{k})-\alpha(1-\alpha)\mu_{k}^{2}.$ (45)
Notice that since $\eta_{k}$ tends to 00, we have that
@@
Hence, by Proposition 5.3, we have that when $n$ and $k$ are large enough,
$f(x_{k+1})-f(\bar{x})\leq f(x_{k})-f(\bar{x})-\frac{1}{2}\alpha(1-\alpha)\left(f(x_{k})-\min\limits_{u\in B_{0}}\tilde{f}_{x_{k}}(u)\right)$ (46)
@@
$f(x_{k+1})-f(\bar{x})\leq\left(1-\frac{1}{2}\alpha(1-\alpha)\frac{\lambda_{0}}{4\lambda_{\max}(\bar{H})\left(\sqrt{\frac{n}{s}}+\mathcal{C}\right)^{2}}\right)\left(f(x_{k})-f(\bar{x})\right),$ (47)
@@
Notice that the rate we obtain corresponds to a high probability estimation of the local convergence rate derived, when $f$ is assumed to be strongly convex, in the Stochastic Subspace Cubic Newton Method [13]. This can be seen in the proof of Proposition 5.5, where the rate $\frac{\lambda_{0}}{4\lambda_{\max}(\bar{H})\left(\sqrt{\frac{n}{s}}+\mathcal{C}\right)^{2}}$ corresponds to a lower bound of $\lambda_{\min}(\bar{H}^{1/2}P_{k}^{\top}(P_{k}\bar{H}P_{k}^{\top})^{-1}P_{k}\bar{H}^{1/2})$, as seen in (34) and (40). More specifically, this corresponds to a high probability lower bound of the parameter $\zeta=\lambda_{\min}[\mathbb{E}(\bar{\Pi})]=\lambda_{\min}[\mathbb{E}(\bar{H}^{1/2}P_{k}^{\top}(P_{k}\bar{H}P_{k}^{\top})^{-1}P_{k}\bar{H}^{1/2})]$ that appears in the local convergence rate in Theorem 6.2 of [13].
@@
Assume that Assumptions 5, 5.1, and 5.1 hold. There exist $k_{0}\in\mathbb{N}$ and $n_{0}\in\mathbb{N}$ such that if $n\geq n_{0}$ and $k\geq k_{0}$, then, for any $m\in\mathbb{N}$,
@@
Proof. This is a direct consequence of Theorem 5.7 where the success probability is obtained by union bound, using (4).
@@
Assume that Assumptions 5, 5.1, and 5.1 hold. There exist $k_{0}\in\mathbb{N}$ and $n_{0}\in\mathbb{N}$ such that if $n\geq n_{0}$ and $k\geq k_{0}$, then,
@@
where $p:=1-6(\exp(-s)+\exp(-\frac{\mathcal{C}_{0}}{4}s))$. Here the expectation is taken with respect to the random variables $P_{1},P_{2},\cdots,P_{k}$.
@@
Proof. By (46) we have that
@@
where $\mathbf{1}_{\mathcal{E}}$ is the indicator function over $\mathcal{E}$. Let us consider the following conditional expectation: $\mathbb{E}\left[\cdot\ |\ P_{1},...,P_{k-1}\right]$. We have that
$\mathbb{E}\left[f(x_{k+1})-f(\bar{x})\ |\ P_{1},...,P_{k-1}\right]\leq f(x_{k})-f(\bar{x})-\frac{1}{2}\alpha(1-\alpha)\mathbb{E}\left[\left(f(x_{k})-\min\limits_{u\in B_{0}}\tilde{f}_{x_{k}}(u)\right)\mathbf{1}_{\mathcal{E}}\ |\ P_{1},...,P_{k-1}\right]$ (48)
@@
=$\mathcal{E}^{\prime}=\left\{f(x_{k})-\min\limits_{u\in B_{0}}\tilde{f}_{x_{k}}(u)\geq\frac{\lambda_{0}}{4\lambda_{\max}(\bar{H})\left(\sqrt{\frac{n}{s}}+\mathcal{C}\right)^{2}}(f(x_{k})-f(\bar{x}))\ |\ x_{k}\right\},$
on this sigma algebra, which holds by probability at least $p=1-6(\exp(-s)+\exp(-\frac{\mathcal{C}_{0}}{4}s))$, by Proposition 5.5. By conditioning the right-hand-side of (48) with respect to this event, we obtain that when $n$ and $k$ are large enough
$\mathbb{E}\left[f(x_{k+1})-f(\bar{x})\ |\ P_{1},\cdots,P_{k-1}\right]$ $\leq f(x_{k})-f(\bar{x})-\frac{1}{2}\alpha(1-\alpha)\mathbb{E}\left[\frac{\lambda_{0}}{4\lambda_{\max}(\bar{H})\left(\sqrt{\frac{n}{s}}+\mathcal{C}\right)^{2}}(f(x_{k})-f(\bar{x}))\mathbf{1}_{\mathcal{E}}\right]p$
@@
Where the first inequality holds as in any case we have that $f(x_{k})-\min\limits_{u\in B_{0}}\tilde{f}_{x_{k}}(u)\geq 0$. By taking the expectation with respect to $P_{1},\cdots,P_{k-1}$ we deduce the corollary.
@@
$\|\nabla f(x_{k})\|\geq\rho\|x_{k}-\bar{x}\|.$ (49)
Notice that Assumption 5.1 is actually stronger than Assumption 5.1$(i)$.
@@
We have, under Assumptions 5 and 5.1, that for $k$ large enough:
@@
Proof. Using a Taylor expansion of $t\mapsto\nabla f(\bar{x}+t(x_{k}-\bar{x}))$ around 00, we have that
@@
By Assumption 5, for any $t\in[0,1]$ we have $\|\nabla^{2}f(\bar{x}+t(x_{k}-\bar{x}))-\bar{H}\|\leq tL_{H}\|x_{k}-\bar{x}\|$. Hence we deduce that
$\|\nabla f(x_{k})\|\leq\|\bar{H}(x_{k}-\bar{x})\|+\|\nabla^{2}f(\bar{x}+t(x_{k}-\bar{x}))-\bar{H}\|\|x_{k}-\bar{x}\|\leq\|\bar{H}(x_{k}-\bar{x})\|+L_{H}\|x_{k}-\bar{x}\|^{2}.$ (51)
Therefore, by (49), we deduce that
[…]
Since $\|x_{k}-\bar{x}\|$ tends to 00, we deduce that for $k$ large enough:
@@
$\|x\|_{\bar{H}}:=x^{\top}\bar{H}x.$ (52)
Notice that by Lemma 5.14, under Assumption 5.1, when $k$ is large enough, $\|\cdot\|_{\bar{H}}$ is a norm for $x_{k}-\bar{x}$ as we have that $\|x_{k}-\bar{x}\|_{\bar{H}}=0$ if and only if $\|x_{k}-\bar{x}\|=0$.
@@
Assume that Assumptions 5, 5.1$(i)$ and 5.1 hold. Then there exists $n_{0}$ such that if $n\geq n_{0}$ and for $k$ large enough:

### v3

@@ 5.1 Local linear convergence
(33) $\bar{H}:=\nabla^{2}f(\bar{x}).$
Also, for $u\in\mathbb{R}^{d}$ small enough, we have by setting $y=x_{k}+P_{k}^{\top}u$ in Eq. 32, that for $k$ large enough such that $x_{k}+P_{k}^{\top}u\in\hat{B}$,
(34) $|f(x_{k}+P_{k}^{\top}u)-f(\bar{x})-\frac{1}{2}(x_{k}-\bar{x})^{\top}\bar{H}(x_{k}-\bar{x})-\frac{1}{2}u^{\top}P_{k}\bar{H}P_{k}^{\top}u-(P_{k}\bar{H}(x_{k}-\bar{x}))^{\top}u|$
@@
holds.
Let $g(u)=\frac{1}{2}u^{\top}P_{k}\bar{H}P_{k}^{\top}u+(P_{k}\bar{H}(x_{k}-\bar{x}))^{\top}u$. By a reasoning similar to that of Proposition 5.1, $g$ is strongly convex with probability $1-6e^{-s}$ and hence is minimized at
(35) $u^{*}=-(P_{k}\bar{H}P_{k}^{\top})^{-1}P_{k}\bar{H}(x_{k}-\bar{x}).$
Notice that as $k$ tends to infinity $\|u^{*}\|$ tends to 0, hence for $k$ large enough we have $x_{k}+P_{k}^{\top}u^{*}\in\hat{B}$ and $u^{*}\in B_{0}$. Plugging Eq. 35 in Eq. 34 yields
@@
where $\Pi=\bar{H}P_{k}^{\top}(P_{k}\bar{H}P_{k}^{\top})^{-1}P_{k}\bar{H}$. Using Eq. 32, we further obtain
(36) $f(x_{k})-f(x_{k}+P_{k}^{\top}u^{*})\geq\frac{1}{2}(x_{k}-\bar{x})^{\top}\Pi(x_{k}-\bar{x})-L_{H}(\|x_{k}-\bar{x}+P_{k}^{\top}u^{*}\|^{3}+\|x_{k}-\bar{x}\|^{3}).$
@@
(37) $=\frac{\lambda_{0}}{2\lambda_{\max}(P_{k}\bar{H}P_{k})}(x_{k}-\bar{x})^{\top}\bar{H}(x_{k}-\bar{x}).$
where the second inequality holds with probability at least $1-2\exp(-\frac{\mathcal{C}_{0}}{4}s)$ (by Lemma 2.1 with $\varepsilon=\frac{1}{2}$), and the third holds as $\lambda_{0}$ is the smallest non-zero eigenvalue of $\bar{H}$. The second equality holds as $\sigma_{\max}(P_{k}\bar{H}^{1/2})^{2}=\lambda_{\max}(P_{k}\bar{H}_{k}P_{k})$. We have therefore proved that
(38) $(\bar{H}^{1/2}(x_{k}-\bar{x}))^{\top}\bar{\Pi}(\bar{H}^{1/2}(x_{k}-\bar{x}))\geq\frac{\lambda_{0}}{2\lambda_{\max}(P_{k}\bar{H}P_{k})}(x_{k}-\bar{x})^{\top}\bar{H}(x_{k}-\bar{x}).$
Hence, by Eq. 36, we have
@@
(39) $-L_{H}(\|x_{k}-\bar{x}+P_{k}^{\top}u^{*}\|^{3}+\|x_{k}-\bar{x}\|^{3}).$
From Eq. 35, we have that $\|x_{k}-\bar{x}+P_{k}^{\top}u^{*}\|=\|(I_{n}-P_{k}^{\top}(P_{k}\bar{H}P_{k}^{\top})^{-1}P_{k}\bar{H})(x_{k}-\bar{x})\|$. Hence
(40) $\|x_{k}-\bar{x}+P_{k}^{\top}u^{*}\|\leq\|I_{n}-P_{k}^{\top}(P_{k}\bar{H}P_{k}^{\top})^{-1}P_{k}\bar{H}\|\|x_{k}-\bar{x}\|.$
@@
$P_{k}\bar{H}P_{k}^{\top}\succeq\frac{n}{8s}\sigma\bar{\lambda}I_{s}.$
Hence, we deduce from Eq. 41 that
(42) $\|I_{n}-P_{k}^{\top}(P_{k}\bar{H}P_{k}^{\top})^{-1}P_{k}\bar{H}\|\leq\frac{\|P_{k}^{\top}\|^{2}\|\bar{H}\|}{\frac{n}{8s}\sigma\bar{\lambda}}.$
Therefore, we deduce by 5.6, Eq. 40 and Eq. 42 for $\beta_{1}>0$ suitably chosen, we have
(43) $f(x_{k})-f(x_{k}+P_{k}^{\top}u^{*})\geq\frac{\lambda_{0}}{4\lambda_{\max}(P_{k}\bar{H}P_{k})}(x_{k}-\bar{x})^{\top}\bar{H}(x_{k}-\bar{x})-\beta_{1}\|x_{k}-\bar{x}\|^{3}.$
By taking $y=x_{k}$ in Eq. 32, we have that
@@
Since $\|x_{k}-\bar{x}\|$ tends to 0 as $k$ tends to infinity and $\rho<3$, we have that for $k$ large enough
@@
where the first inequality holds as, by Eq. 35, $u^{*}\in B_{0}$ for $k$ large enough. The probability bound in the statement of the theorem is obtained by using Eq. 4: in the whole proof we only use Lemma 2.1 with $\varepsilon=\frac{1}{2}$, which holds with probability at least $1-2\exp(-\frac{\mathcal{C}_{0}}{4}s)$, and the inequalities Eq. 29 which hold with probability at least $1-6\exp(-s)$. We also factorize the expression, using that $1-2\exp(-\frac{\mathcal{C}_{0}}{4}s)>1-6\exp(-\frac{\mathcal{C}_{0}}{4}s)$. We end the proof by noticing that $\lambda_{\max}(P_{k}\bar{H}P_{k})\leq\lambda_{\max}(\bar{H})\sigma_{\max}(P_{k})^{2}$, hence by the first equation of Eq. 29
(44) $\lambda_{\max}(P_{k}\bar{H}P_{k})\leq\lambda_{\max}(\bar{H})\left(\sqrt{\frac{n}{s}}+\mathcal{C}\right)^{2}.$
@@
Assume that Section 5, Section 5.1 and Section 5.1 hold. There exist $0<\kappa<1$, $k_{0}\in\mathbb{N}$, such that if $k\geq k_{0}$, then
@@
with $d_{k}=P_{k}^{\top}u_{k}$ and the update rule $t_{k}\leftarrow\beta t_{k}$ for $0<\alpha<1$ and $0<\beta<1$. We recall that
(45) $u_{k}=-(P_{k}H_{k}P_{k}^{\top}+\eta_{k}I_{s})^{-1}P_{k}g_{k},$
@@
(46) $f(x_{k}+t_{k}P_{k}^{\top}u_{k})=f(x_{k})+t_{k}(P_{k}g_{k})^{\top}u_{k}+\frac{t_{k}^{2}}{2}u_{k}^{\top}P_{k}\nabla^{2}f(x_{k}^{*})P_{k}^{\top}u_{k}.$
Notice that $\nabla^{2}f$ is Lipschitz continuous (by Section 5). Furthermore, by Proposition 5.1, for $k$ large enough, $P_{k}{H}_{k}P_{k}^{\top}$ is positive definite with probability at least $1-6\exp(-s)$ as the sequence $\{x_{k}\}$ converges to $\bar{x}$. Hence, for $k$ large enough
@@
holds with probability at least $1-6(\exp(-s)+\exp(-\frac{\mathcal{C}_{0}}{4}s))$. By Eq. 46, we deduce that for $k$ large enough:
$f(x_{k}+t_{k}P_{k}^{\top}u_{k})$ $\leq f(x_{k})+t_{k}(P_{k}g_{k})^{\top}u_{k}+2\frac{t_{k}^{2}}{2}u_{k}^{\top}P_{k}{H}_{k}P_{k}^{\top}u_{k}$
$\leq f(x_{k})+t_{k}(P_{k}g_{k})^{\top}u_{k}+{t_{k}^{2}}u_{k}^{\top}(P_{k}{H}_{k}P_{k}^{\top}+\eta_{k}I_{s})u_{k},$
where the second inequality holds as $\eta_{k}\geq 0$. Let
(47) $\mu_{k}^{2}:=-g_{k}^{\top}d_{k}=(P_{k}g_{k})^{\top}(P_{k}H_{k}P_{k}^{\top}+\eta_{k}I_{s})^{-1}(P_{k}g_{k}).$
Since $(P_{k}g_{k})^{\top}u_{k}=g_{k}^{\top}(P_{k}^{\top}u_{k})=-\mu_{k}^{2}$, and by definition of $u_{k}$ in Eq. 45, we can write
(48) $f(x_{k}+t_{k}P_{k}^{\top}u_{k})\leq f(x_{k})-t_{k}\mu_{k}^{2}+{t_{k}^{2}}u_{k}^{\top}(P_{k}{H}_{k}P_{k}^{\top}+\eta_{k}I_{s})u_{k}=f(x_{k})-t_{k}\mu_{k}^{2}+{t_{k}^{2}}\mu_{k}^{2}.$
@@
for such $t_{k}$. Therefore, the backtracking line search stops with some $t_{k}\geq 1-\alpha$, and we have
(49) $f(x_{k+1})\leq f(x_{k})-\alpha(1-\alpha)\mu_{k}^{2}.$
Notice that since $\eta_{k}$ tends to 0, we have that
@@
Hence, by Proposition 5.3, we have that when $k$ is large enough,
(50) $f(x_{k+1})-f(\bar{x})\leq f(x_{k})-f(\bar{x})-\frac{1}{2}\alpha(1-\alpha)\left(f(x_{k})-\min\limits_{u\in B_{0}}\tilde{f}_{x_{k}}(u)\right)$
@@
(51) $f(x_{k+1})-f(\bar{x})\leq\left(1-\frac{1}{2}\alpha(1-\alpha)\frac{\lambda_{0}}{4\lambda_{\max}(\bar{H})\left(\sqrt{\frac{n}{s}}+\mathcal{C}\right)^{2}}\right)\left(f(x_{k})-f(\bar{x})\right),$
@@
Notice that the rate we obtain corresponds to a high probability estimation of the local convergence rate derived, when $f$ is assumed to be strongly convex, in the stochastic subspace cubic Newton method [22]. This can be seen in the proof of Proposition 5.5, where the rate $\frac{\lambda_{0}}{4\lambda_{\max}(\bar{H})\left(\sqrt{\frac{n}{s}}+\mathcal{C}\right)^{2}}$ corresponds to a lower bound of $\lambda_{\min}(\bar{H}^{1/2}P_{k}^{\top}(P_{k}\bar{H}P_{k}^{\top})^{-1}P_{k}\bar{H}^{1/2})$, as seen in Eq. 38 and Eq. 44. More specifically, this corresponds to a high probability lower bound of the parameter $\zeta=\lambda_{\min}[\mathbb{E}(\bar{\Pi})]=\lambda_{\min}[\mathbb{E}(\bar{H}^{1/2}P_{k}^{\top}(P_{k}\bar{H}P_{k}^{\top})^{-1}P_{k}\bar{H}^{1/2})]$ that appears in the local convergence rate in Theorem 6.2 of [22].
@@
Assume that Section 5, Section 5.1 and Section 5.1 hold. There exist $k_{0}\in\mathbb{N}$ such that if $k\geq k_{0}$, then, for any $m\in\mathbb{N}$,
@@
Proof. This is a direct consequence of Theorem 5.7 where the success probability is obtained by union bound, using Eq. 4.
@@
Assume that Section 5, Section 5.1 and Section 5.1 hold. There exist $k_{0}\in\mathbb{N}$ such that if $k\geq k_{0}$, then,
@@
where $p:=1-6(\exp(-s)+\exp(-\frac{\mathcal{C}_{0}}{4}s))$. Here the expectation is taken with respect to the random variables $P_{0},P_{1},P_{2},\cdots,P_{k}$.
@@
Proof. By Eq. 50 we have that
@@
where $\mathbf{1}_{\mathcal{E}}$ is the indicator function over $\mathcal{E}$. Let us consider the following conditional expectation: $\mathbb{E}\left[\cdot\ |\ P_{0},...,P_{k-1}\right]$. We have that
(52) $\mathbb{E}\left[f(x_{k+1})-f(\bar{x})\ |\ P_{0},...,P_{k-1}\right]\leq f(x_{k})-f(\bar{x})-\frac{1}{2}\alpha(1-\alpha)\mathbb{E}\left[\left(f(x_{k})-\min\limits_{u\in B_{0}}\tilde{f}_{x_{k}}(u)\right)\mathbf{1}_{\mathcal{E}}\ |\ P_{0},...,P_{k-1}\right]$
@@
$\mathcal{E}^{\prime}=\left\{f(x_{k})-\min\limits_{u\in B_{0}}\tilde{f}_{x_{k}}(u)\geq\frac{\lambda_{0}}{4\lambda_{\max}(\bar{H})\left(\sqrt{\frac{n}{s}}+\mathcal{C}\right)^{2}}(f(x_{k})-f(\bar{x}))\ |\ x_{k}\right\},$
on this sigma algebra, which holds by probability at least $p=1-6(\exp(-s)+\exp(-\frac{\mathcal{C}_{0}}{4}s))$, by Proposition 5.5. By conditioning the right-hand-side of Eq. 52 with respect to this event, we obtain that when $k$ is large enough
$\mathbb{E}\left[f(x_{k+1})-f(\bar{x})\ |\ P_{0},\cdots,P_{k-1}\right]$ $\leq f(x_{k})-f(\bar{x})-\frac{1}{2}\alpha(1-\alpha)\mathbb{E}\left[\frac{\lambda_{0}}{4\lambda_{\max}(\bar{H})\left(\sqrt{\frac{n}{s}}+\mathcal{C}\right)^{2}}(f(x_{k})-f(\bar{x}))\mathbf{1}_{\mathcal{E}}\right]p$
@@
Where the first inequality holds as in any case we have that $f(x_{k})-\min\limits_{u\in B_{0}}\tilde{f}_{x_{k}}(u)\geq 0$. By taking the expectation with respect to $P_{0},\cdots,P_{k-1}$ we deduce the corollary.
@@
(53) $\|\nabla f(x_{k})\|\geq\rho\|x_{k}-\bar{x}\|.$
Notice that Section 5.1 is actually stronger than Section 5.1(ii)(ii).
@@
We have, under Section 5 and Section 5.1, that for $k$ large enough:
@@
Proof. Using a Taylor expansion of $t\mapsto\nabla f(\bar{x}+t(x_{k}-\bar{x}))$ around 0, we have that
(54) $\nabla f(x_{k})=\nabla f(\bar{x})+\int_{0}^{1}\nabla^{2}f(\bar{x}+t(x_{k}-\bar{x}))(x_{k}-\bar{x})dt=\int_{0}^{1}\nabla^{2}f(\bar{x}+t(x_{k}-\bar{x}))(x_{k}-\bar{x})dt.$
By Section 5, for any $t\in[0,1]$ we have $\|\nabla^{2}f(\bar{x}+t(x_{k}-\bar{x}))-\bar{H}\|\leq tL_{H}\|x_{k}-\bar{x}\|$. Hence we deduce that
(55) $\|\nabla f(x_{k})\|\leq\|\bar{H}(x_{k}-\bar{x})\|+\|\nabla^{2}f(\bar{x}+t(x_{k}-\bar{x}))-\bar{H}\|\|x_{k}-\bar{x}\|\leq\|\bar{H}(x_{k}-\bar{x})\|+L_{H}\|x_{k}-\bar{x}\|^{2}.$
Therefore, by Eq. 53, we deduce that
@@
Since $\|x_{k}-\bar{x}\|$ tends to 0, we deduce that for $k$ large enough:
@@
(57) $\|x\|^{2}_{\bar{H}}:=x^{\top}\bar{H}x.$
Notice that by Lemma 5.14, under Section 5.1, when $k$ is large enough, $\|\cdot\|_{\bar{H}}$ is a norm for $x_{k}-\bar{x}$ as we have that $\|x_{k}-\bar{x}\|_{\bar{H}}=0$ if and only if $\|x_{k}-\bar{x}\|=0$.
@@
Assume that Section 5, Section 5.1, Section 5.1(i)(i) and Section 5.1 hold. Then for $k$ large enough: